OpenStudy (anonymous):

solve for x: 7x^2 - 3x - 4 = 0

OpenStudy (anonymous):

you can use the formula :)

OpenStudy (anonymous):

can you show me how it's done step by step? i forget

OpenStudy (anonymous):

when you have a equation like this: ax^2 + bx + c = 0 you can use this formula: \[x _{1,2} = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

but you can solve it by the laws of mathematics too :)

OpenStudy (anonymous):

i prefer to show you how to do that by the laws of mathematics

OpenStudy (anonymous):

@sexychick

OpenStudy (anonymous):

7x^2-3x-4=0 (7x+4)(x-1)=0 7x+4=0 or x-1=0 7x=-4 0r x=1 x=-4/7 or x=1 solutions are -4/7 and 1 hope that helps

OpenStudy (anonymous):

@melbe , your explanation was nice but please explain \( \sf \large \color{green}{Step} \; \color{red}{by} \; \color{green}{step} \)